OpenStudy (anonymous):

Divide (8a^-8ab-6b^) by (2a - 3b) guys... again.. i need ur help...

OpenStudy (mathstudent55):

First, write it as a fraction: \( \dfrac{8a^{2}-8ab-6b^{2}}{2a - 3b} \) Do you mean that division? You don't have exponents on a and b.

OpenStudy (anonymous):

wow... yes..

OpenStudy (anonymous):

after?

OpenStudy (mathstudent55):

Now you need to factor the numerator if possible and the denominator if possible.

OpenStudy (anonymous):

so whats ur final answer in the end?

OpenStudy (mathstudent55):

\( \dfrac{8a^{2}-8ab-6b^{2}}{2a - 3b} \) Factor the numerator: \( =\dfrac{2(4a^{2}-4ab-3b^{2})}{2a - 3b} \) \( =\dfrac{2(2a-3b)(2a + b)}{2a - 3b} \) Now cancel out the common term in the numerator and denominator.

OpenStudy (anonymous):

thank u so much @mathstudent55 ..

OpenStudy (mathstudent55):

yw

OpenStudy (anonymous):

2 (2a+b) ?? thats the answer?

OpenStudy (mathstudent55):

yes